1. 3 points) What is the current through, and voltage drop across, each resistor in the following circuit? 3.0 9.0v (3 points) A beam of protons travels at 3.5 x 10 m/s through a uniform magnetic field of 5.7 x 102 T. The beam is perpendicular to the field. How strong is the magnetic force acting on each proton? The proton charge is +1.6 x10-19 C.
Answer #1

v = 3.5*10^6 m/s

q = 1.6*10^-19 C

B = 5.7*10^-2 T

theta = 90 degree

now

Force = F = qvB*sin(theta) = 1.6*10^-19*3.5*10^6*5.7*10^-2*sin90 = 3.2*10^-14 N